# GPT4Rec: Graph Prompt Tuning for Streaming Recommendation

### v1

@@ 3.1. Definitions and Formulations
Definition 1. Streaming Recommendation. Massive user-item interaction data $\tilde{D}$ streams into industrial recommender system continuously. For convenience (Wang et al., 2020, 2022b; Zhang et al., 2023c), the continuous data stream is split into consecutive data segments $D_{1},...,D_{t},...,D_{T}$ with the same time span. At each time segment $t$ , the model needs to optimize the recommendation performance on $D_{t}$ with the knowledge inherited from $D_{1},D_{2},...,D_{t-1}$. The recommendation performance is evaluated along the whole timeline.
Definition 2. Streaming Graph. A streaming graph is represented as a sequence of graphs $G=(G_{1},G_{2},...,G_{t},...G_{T})$, where $G_{t}=G_{t-1}+\Delta G_{t}$. $G_{t}=(A_{t},X_{t})$ is an attributed graph at time $t$, where $A_{t}$ and $X_{t}$ are the adjacency matrix and node features of $G_{t}$ , respectively. $\Delta G_{t}=(\Delta A_{t},\Delta X_{t})$ is the changes of graph structures and node attributes at $t$ The changes contain newly added nodes and newly built connections between different nodes.
Definition 3. Continual Graph Learning for Streaming Graph. Given a streaming graph $G=(G_{1},G_{2},...,G_{t},...G_{T})$, the goal of continual graph learning (CGL) is to learn $\Delta G_{t}(D_{t})$ sequentially while transferring historical knowledge to new graph segments effectively. Mathematically, the goal of CGL for streaming graph is to find the optimal GNN structure $S_{t}$ and parameters $W_{t}$ at each segment $t$ such that:
@@
where $S_{t},W_{t}\in(\mathcal{S},\mathcal{W})$. $\mathcal{L}_{t}(S_{t},W_{t},\Delta G_{t})$ is the loss function of current task defined on $\Delta G_{t}$ The $\mathcal{S}$ and $\mathcal{W}$ are corresponding search spaces, respectively.
Since the user-item interaction data is actually a bipartite graph, the continual learning task for streaming recommendation is essentially the continual graph learning for streaming graph. For each segment $t$, the GNN structure $S_{t}$ and parameters $W_{t}$ need to be adjusted and refined simultaneously to achieve a satisfying recommendation performance. We use the Bayesian Personalized Ranking (BPR) (Rendle et al., 2012) loss as the loss function in this work, because it is effective and has broad applicability in top-K recommendations.
@@ 4. Methodology
In this section, we introduce our GPT4Rec method towards continual graph learning for streaming recommendation. We first model the temporal user preference to capture the preference shifts between segments, which will be utilized for the user embedding initialization of training. Then, we successively use two operations, historical graph convolution pruning and refining as well as graph convolution expanding and pruning (shown in Figure 2), in such a model isolation way to obtain the best structure and optimal parameters of the graph convolution.
@@ 4.1. Disentangling Strategy for Complex Graphs
where ${Linear}_{i}$ is the linear transformation of the $i$-th view. The linear transformations are capable of separating these views while maintaining the overall integrity of the graph. This means the model can isolate and focus on specific aspects without losing sight of the graph’s interconnected nature. This disentanglement allows the model to explore and identify distinct patterns and relationships in the data. For example, one view might capture user-to-item interactions, another might focus on user-to-user relationships, and yet another could delve into item-to-item similarities. With each view focusing on a particular aspect of the graph, the model can adapt more precisely to changes or updates in the data.
@@ 4.2. Prompt Design for Adaptive Learning
(8) $att_{ij,k}=\frac{exp(a(\tilde{x}_{i}||\tilde{x}_{j})q_{k})}{\Sigma_{r}exp(a(\tilde{x}_{i}||\tilde{x}_{j})q_{r})},$
where $x_{i}$ and $x_{j}$ are adjacent nodes, $a$ is the linear mapping.
@@ 4.3. Aggregation of Disentangled Representations
(9) $\hat{x}_{i}=Atten(p(x_{i}),[\tilde{x}_{i,1},...,\tilde{x}_{i,n}])=x_{i}+\Sigma_{j}\gamma_{i,j}\tilde{x}_{i,j},$
(10) $\gamma_{i,j}=\frac{exp(p(x_{i})\tilde{x}_{i,j})}{\Sigma_{r}exp(p(x_{i}),\tilde{x}_{i,r})},$
where $p()$ is a linear transformation function. However, this approach may not fully account for the evolving nature of user-item interactions, especially with the introduction of new data. As the graph’s structure changes, the relevance of different views and their interrelations can shift, potentially rendering the existing fixed attention weights less effective.
@@
To enhance the model’s efficiency and adaptability in the face of these dynamic changes, GPT4Rec incorporates Cross-View-Level Prompts for dynamic adaptation. This approach centers on updating a small set of ‘codebook’ prompts $J=[j_{1},...,j_{n}]$, rather than relearning the entire model’s parameters. These prompts serve as dynamic modifiers to the attention mechanism, allowing the model to adapt its focus efficiently:
(11) $\hat{x}_{i}=Atten(prompt(x_{i}),[\tilde{x}_{i,1},...,\tilde{x}_{i,n}])=x_{i}+\Sigma_{j}\epsilon_{i,j}\tilde{x}_{i,j},$
(12) $\epsilon{i,j}=\frac{exp(prompt(x_{i})\tilde{x}_{i,j})}{\Sigma_{r}exp(prompt(x_{i}),\tilde{x}_{i,r})}=\frac{exp(p(x_{i}+j_{i})\tilde{x}_{i,j})}{\Sigma_{r}exp(p(x_{i}+j_{i}),\tilde{x}_{i,r})},$
@@ 5.1. Experiment Settings
We apply grid search to find the optimal hyper-parameters for each model. The ranges of hyper-parameters are $\{32,64,96,128\}$ for size $L$ of node-level prompts $P$, the size $K$ of structure-level prompts $Q$ and the size $N$ of cross-view-level prompts $J$. The range of disentangled view number is $\{2,4,8,16\}$. Adam optimizer (Kingma and Ba, 2014) is employed to minimize the training loss. Other parameters are tuned on the validation dataset and we save the checkpoint with the best validation performance as the final model. We use the same evaluation metrics Recall@K (abbreviated as R@K) and NDCG@K (abbreviated as N@K) following previous studies (Xu et al., 2020; Wang et al., 2021; He et al., 2023). All models are run five times with different random seeds and reported the average on a single NVIDIA GeForce RTX 3090 GPU.
@@ 5.3. Method Robustness Analysis
To figure out whether our method is robust to different datasets and GCN-based recommendation models, we conduct the experiments with NGCF and LightGCN as the base GCN models on both Taobao2014 and Netflix datasets. The corresponding results are shown in Tables 3 and Table 4. For GCN models NGCF and LightGCN, the improvements of our methods on Taobao2014 are both significant. GPT4Rec achieves the state-of-art recommendation performance. This also shows the performance potential of GPT4Rec on different kinds of base GCN models. As for the dataset Netflix, GPT4Rec improves the Recall@20 by 6.71% and NDCG@20 by 5.91% over Finetune when choosing NGCF as the GCN model. Similar improvements can also be observed when taking LightGCN as the base GCN model. Such observations demonstrate the robustness of our methods to different datasets and GNN backbones.
@@ 5.4. Efficiency
Figure 2 presents the average training time of GPT4Rec per epoch compared with baselines. From the results, we observe that GPT4Rec not only aligns closely with the efficiency of the Finetune approach but also surpasses several other advanced models in terms of training speed. The high efficiency of GPT4Rec can be primarily attributed to the utilization of lightweight graph prompts. These prompts, despite their minimal computational footprint, play a crucial role in seamlessly integrating new data into the model. By leveraging these compact yet effective prompts, GPT4Rec bypasses the need for extensive retraining or large-scale parameter adjustments typically required by other models. This light-weight approach ensures rapid adaptability to new information, significantly reducing the computational overhead and training time.
@@ 5.5. Ablation Study
  • The process of decomposing the graph into distinct views is a fundamental aspect of the GPT4Rec model. This disentanglement facilitates the model’s understanding by allowing it to separately process diverse interaction dynamics between users and items. The results suggest that disentangling the graph into multiple views substantially improves the model’s capability to accurately represent and interpret the multifaceted relationship dynamics. Specifically, the disentangled views enable the model to address different aspects of the graph’s structure in isolation, thus enhancing the quality of its recommendations by providing a more precise understanding of the interaction patterns.
@@
  • Finally, the view aggregation component is scrutinized. This aspect is crucial for reintegrating the disentangled views into a cohesive model output. The ablation study shows that effective aggregation is key to ensuring that the insights gained from the separate views are synergistically combined, leading to a more comprehensive understanding of the graph data.
@@ 5.6. Hyperparameter Study
We first study the impact of varying the size of node-level prompts $P$ with size $L$, structure-level prompts $Q$ with size $K$, and cross-view-level prompts $J$ with size $N$ on the Tb2014 dataset. The results are shown in Figure 4. We find that initially increasing the size of these prompts generally enhances the performance, indicating that Larger prompt sizes allow for a richer representation of complex user-item interactions and relationships, providing the model with a more diverse set of ”hints” to interpret and integrate new information effectively. When the prompt size is large enough, we notice that further increasing the prompt size only brings marginal benefits. Moreover, excessively large prompts introduce additional computational overhead, which may not be justifiable given the marginal gains in performance.
@@
We explore the influence of the number of disentangled views on the Tb2014 dataset. The results are shown in Figure 5. We find that an increase in the number of disentangled views generally corresponds to improved performance, primarily due to each view providing a distinct lens through which the model can perceive and process various aspects of user-item interactions. This multi-view approach enables a richer, more layered understanding of the data, as each view contributes unique insights into different facets of the graph, such as varying user preferences or item characteristics.

### v2

@@ 3.1. Definitions and Formulations
Definition 1. Streaming Recommendation. Industrial recommender systems handle a continuous influx of user-item interaction data, denoted as $\tilde{D}$. This stream is divided into sequential segments $D_{1},...,D_{t},...,D_{T}$, each covering an equal time period (Wang et al., 2022b; Zhang et al., 2023c). In each time segment $t$, the model optimizes recommendation performance on $D_{t}$ by incorporating knowledge from previous segments $D_{1},D_{2},...,D_{t-1}$. The performance is evaluated over the entire timeline.
Definition 2. Streaming Graph. A streaming graph is defined as a series of graphs $G=(G_{1},G_{2},...,G_{t},...G_{T})$, where each graph $G_{t}$ evolves from its predecessor by incorporating changes $\Delta G_{t}$ such that $G_{t}=G_{t-1}+\Delta G_{t}$. At time $t$, the attributed graph $G_{t}=(A_{t},X_{t})$ consists of an adjacency matrix $A_{t}$ and node feature matrix $X_{t}$. The incremental changes $\Delta G_{t}=(\Delta A_{t},\Delta X_{t})$ reflect modifications in graph structure and node attributes, including the addition/deletion of nodes and edges.
Definition 3. Continual Graph Learning (CGL) for Streaming Graph. Given a streaming graph $G=(G_{1},G_{2},...,G_{t},...G_{T})$, the objective of CGL is to sequentially learn the updates $\Delta G_{t}(D_{t})$ while effectively transferring knowledge from historical data to new graph segments. Formally, CGL aims to learn the optimal graph structure $S_{t}$ and parameters $W_{t}$ for each segment $t$, formulated as:
@@
where $S_{t}$ and $W_{t}$ are elements of the search spaces $\mathcal{S}$ and $\mathcal{W}$, respectively. The function $\mathcal{L}{t}(S{t},W_{t},\Delta G_{t})$ denotes the loss for the current segment, evaluated on $\Delta G_{t}$.
@@ 4. Methodology
In this section, we introduce our GPT4Rec method towards continual graph learning for streaming recommendation. As shown in Figure 1, we first disentangle the complex user-item interaction graphs into multiple views, capturing specific interaction aspects. We then design node-level prompts to focus on individual node attributes and structure-level prompts to address broader connectivity patterns. Afterwards, we propose the cross-view-level prompts to aggregate information from these views, integrating new knowledge efficiently and adaptively.
@@ 4.1. Disentangling Strategy for Complex Graphs
where ${Linear}_{i}$ is the linear transformation of the $i$-th view, $i\in\{1,2,...,N\}$. The linear transformations are capable of separating these views while maintaining the overall integrity of the graph. This means the model can isolate and focus on specific aspects without losing sight of the graph’s interconnected nature. This disentanglement allows the model to explore and identify distinct patterns and relationships in the data. For example, one view might capture user-to-item interactions, another might focus on user-to-user relationships, and yet another could delve into item-to-item similarities. With each view focusing on a particular aspect of the graph, the model can adapt more precisely to changes or updates in the data.
@@ 4.2. Prompt Design for Adaptive Learning
(8) $att_{ij,k}=\frac{exp({Linear}(\tilde{x}_{i}||\tilde{x}_{j})q_{k})}{\Sigma_{r}exp({Linear}(\tilde{x}_{i}||\tilde{x}_{j})q_{r})},$
where $x_{i}$ and $x_{j}$ are adjacent nodes.
@@ 4.3. Aggregation of Disentangled Representations
(9) $\hat{x}_{i}=Atten({Linear}(x_{i}),[\tilde{x}_{i,1},...,\tilde{x}_{i,n}])=x_{i}+\Sigma_{j}\gamma_{i,j}\tilde{x}_{i,j},$
(10) $\gamma_{i,j}=\frac{exp({Linear}(x_{i})\tilde{x}_{i,j})}{\Sigma_{r}exp({Linear}(x_{i})\tilde{x}_{i,r})},$
However, this approach may not fully account for the evolving nature of user-item interactions, especially with the introduction of new data. As the graph’s structure changes, the relevance of different views and their interrelations can shift, potentially rendering the existing fixed attention weights less effective.
@@
To enhance the model’s efficiency and adaptability in the face of these dynamic changes, GPT4Rec incorporates Cross-View-Level Prompts for dynamic adaptation. This approach centers on updating a small set of ‘codebook’ prompts $Z=[z_{1},...,z_{N}]$, rather than relearning the entire model’s parameters. These prompts serve as dynamic modifiers to the attention mechanism, allowing the model to adapt its focus efficiently:
(11) $\hat{x}_{i}=att(prompt(x_{i}),[\tilde{x}_{i,1},...,\tilde{x}_{i,N}])=x_{i}+\Sigma_{j}^{N}\epsilon_{(i,z,j)}\tilde{x}_{i,j},$
(12) $\epsilon{(i,z,j)}=\frac{exp(prompt(x_{i})\tilde{x}_{i,j})}{\Sigma_{r}exp(prompt(x_{i})\tilde{x}_{i,r})}=\frac{exp({Linear}(x_{i}+z_{i})\tilde{x}_{i,j})}{\Sigma_{r}exp({Linear}(x_{i}+z_{i})\tilde{x}_{i,r})},$
@@ 5.1. Experiment Settings
We apply grid search to find the optimal hyper-parameters for each model. The ranges of hyper-parameters are $\{32,64,96,128\}$ for size $L$ of node-level prompts $P$, the size $K$ of structure-level prompts $Q$ and the size $N$ of cross-view-level prompts $Z$. The range of disentangled view number is $\{2,4,8,16\}$. Adam optimizer (Kingma and Ba, 2014) is employed to minimize the training loss. Other parameters are tuned on the validation dataset and we save the checkpoint with the best validation performance as the final model. We use the same evaluation metrics Recall@K (abbreviated as R@K) and NDCG@K (abbreviated as N@K) following previous studies (Xu et al., 2020; Wang et al., 2021; He et al., 2023). All models are run five times with different random seeds and reported the average on a single NVIDIA GeForce RTX 3090 GPU.
@@ 5.3. Method Robustness Analysis
To evaluate the robustness of our method across various datasets and different GNN backbones, we use different GNN models as the backbones of GPT4Rec, and evaluate on the Netflix and Taobao2014 datasets. The results can be found in Tables 3 and 4.
We observe that our method consistently outperform baselines on both datasets. Moreover, we find that our GPT4Rec can be combined with different GNN backbones properly. For example, on Taobao2014 dataset, when using NGCF as the backbone GNN, GPT4Rec outperforms the Finetune baseline by 6.71% and 5.91% in terms of Recall@20 and NDCG@20, respectively. These results underscore the robustness and adaptability of GPT4Rec to different datasets and GNN backbones, indicating its potential for broad applicability in diverse recommendation scenarios.
@@ 5.4. Efficiency
Figure 2 illustrates the average training time per epoch of GPT4Rec in comparison with various baseline models. The results reveal that GPT4Rec not only matches the efficiency of the Finetune approach but also outperforms several other advanced models in terms of training speed. This high efficiency can be primarily attributed to the use of lightweight graph prompts. These prompts, despite their minimal computational demands, play a crucial role in seamlessly integrating new data into the model. By leveraging these compact yet powerful prompts, GPT4Rec avoids the need for extensive retraining or large-scale parameter adjustments that other models typically require. This streamlined approach ensures rapid adaptability to new information, significantly reducing computational overhead and training time, thereby enhancing overall efficiency.
@@ 5.5. Ablation Study
  • The view aggregation component is scrutinized. This aspect is crucial for reintegrating the disentangled views into a cohesive model output. The ablation study shows that effective aggregation is key to ensuring that the insights gained from the separate views are synergistically combined, leading to a more comprehensive understanding of the graph data.
  • Finally, by comparing GPT4Rec’s performance against baseline models that lack these features, it becomes evident that each component plays a vital role. The integration of view disentanglement, node-level prompts, structure-level prompts, and effective view aggregation sets GPT4Rec apart from its counterparts. This comprehensive design enables GPT4Rec to not only accurately capture the evolving nature of user-item interactions but also to do so in a way that is both efficient and scalable.
@@ 5.6. Hyperparameter Study
We first study the impact of varying the size of node-level prompts $P$ with size $L$, structure-level prompts $Q$ with size $K$, and cross-view-level prompts $Z$ with size $N$ on the Tb2014 dataset. The results are shown in Figure 4.
We find that initially increasing the size of these prompts generally enhances performance. Larger prompt sizes allow for a richer representation of complex user-item interactions and relationships, providing the model with a more diverse set of ”hints” to interpret and integrate new information effectively. This improvement can be attributed to the model’s enhanced ability to capture and utilize the subtleties of user preferences and item characteristics. However, when the prompt size reaches a certain threshold, further increases only bring marginal benefits. Beyond this point, the additional capacity of the prompts does not significantly improve the model’s interpretative power. Instead, the gains in performance become incremental and less impactful. Moreover, excessively large prompts introduce additional computational overhead, which may not be justifiable given the marginal gains in performance.
@@
We explore the influence of the number of disentangled views on the Tb2014 dataset. The results are shown in Figure 5. We find that the process of decomposing the graph into distinct views is a fundamental aspect of the GPT4Rec model. This disentanglement facilitates the model’s understanding by allowing it to separately process diverse interaction dynamics between users and items. The results suggest that disentangling the graph into multiple views substantially improves the model’s capability to accurately represent and interpret the multifaceted relationship dynamics. Specifically, the disentangled views enable the model to address different aspects of the graph’s structure in isolation, thus enhancing the quality of its recommendations by providing a more precise understanding of the interaction patterns.
In addition, an increase in the number of disentangled views generally corresponds to improved performance, primarily due to each view providing a distinct lens through which the model can perceive and process various aspects of user-item interactions. This multi-view approach enables a richer, more layered understanding of the data, as each view contributes unique insights into different facets of the graph, such as varying user preferences or item characteristics.